OpenStudy (anonymous):

Solve for x: -2(x - 3) = 10

OpenStudy (anonymous):

x=-2

OpenStudy (anonymous):

-2(x-3) = 10 -2x+6=10 -6 -6 -2x=4 ------- 4 4 x=-2 -2(-2-3)=10